OpenStudy (anonymous):

7x7+(6-2)=n 7x7+(6-2)=n @Mathematics

OpenStudy (anonymous):

7x7+(6-2)=n 7x7+(4)=n 49x4=n 196 = n

OpenStudy (anonymous):

so the answer is

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

196. It's an order of operation problem

OpenStudy (anonymous):

thats not right the choices r 60 or 77or 55 or 53

OpenStudy (anonymous):

Woops, it should be 53

OpenStudy (anonymous):

so, it would PEMDAS

OpenStudy (anonymous):

o k

OpenStudy (anonymous):

7x7+(6-2)=n 7x7+(4)=n 49+4=n 53 = n

OpenStudy (anonymous):

be using PEMDAS rite?

OpenStudy (anonymous):

medal given

OpenStudy (anonymous):

that's exactly what you're using :)

OpenStudy (anonymous):

oh k thx